# Learning Constrained Markov Decision Processes With Non-stationary Rewards and Constraints

### v1

@@ F.1 Lagrangian formulation of the constrained optimization problem
$\min_{\lVert\lambda\rVert_{1}\in[0,L/\rho]}\max_{q\in\Delta(P)}\mathcal{Q}(\lambda,q)=\min_{\lambda\in\mathbb{R}^{m}{}^{+}}\max_{q\in\Delta(P)}\mathcal{Q}(\lambda,q),$
@@
$\min_{\lambda\in\mathbb{R}^{m}{}^{+}}\max_{q\in\Delta(P)}\mathcal{Q}(\lambda,q)$ $=\min\left\{\min_{\lVert\lambda\rVert_{1}\in[0,L/\rho]}\max_{q\in\Delta(P)}\mathcal{Q}(\lambda,q),\min_{\lVert\lambda\rVert_{1}\geq L/\rho}\max_{q\in\Delta(P)}\mathcal{Q}(\lambda,q)\right\}$
@@
$\text{OPT}_{r,G,\alpha}$ $=\max_{q\in\Delta(P)}\min_{\lambda\in\mathbb{R}^{m}{}^{+}}\mathcal{Q}(\lambda,q)$
@@ F.2 Preliminary results
For any $\delta\in(0,1)$ and properly setting the learning rate $\eta$ such that $\eta\leq\frac{\rho}{2Lm(Lm+1)}$, Algorithm 3 attains, with probability at least $1-2\delta$:
@@
$\leq\mathcal{O}\left(\frac{M\ln T}{\eta}+\eta~{}{m^{4}L^{4}TM}+\gamma TLM+L\sqrt{T\ln\left(\frac{1}{\delta}\right)}+\frac{L}{\gamma}\ln\left(\frac{1}{\delta}\right)\right).$
@@
$\ell_{t,j}:=\frac{\mathbb{I}(j_{t}=j)}{w_{t,j}+\gamma}\left(\sum_{(x_{k}^{t},a_{k}^{t})}(1-r_{t}(x_{k}^{t},a_{k}^{t}))+\frac{Lm+1}{\rho}\sum_{i\in[m]}\left[\widehat{g}_{t,i}^{\ j\top}\widehat{q}_{t}^{\ j}-\alpha_{i}\right]^{+}\right).$
Since it holds that $\eta w_{t,j}\lvert\ell_{t,j}\rvert\leq\frac{\eta(L\rho+L^{2}m^{2}+Lm)}{\rho}\leq\frac{1}{2}$ for all $j\in[M]$, for all $t\in[T]$ as long as $\eta\leq\frac{\rho}{2(L\rho+L^{2}m^{2}+Lm)}\leq\frac{\rho}{2(L^{2}m^{2}+Lm)}$, if $\eta\leq\frac{\rho}{2Lm(Lm+1)}$, Algorithm 3 attains, by Lemma 17 :
@@
$\leq\frac{M\ln T}{\eta}+\eta\frac{TM(L\rho+L^{2}m^{2}+Lm)^{2}}{\rho^{2}}+\frac{Lm+1}{\rho}\widehat{V}_{T,j^{*}},$ (18)
where we used that $\left(\sum_{(x_{k}^{t},a_{k}^{t})}(1-r_{t}(x_{k}^{t},a_{k}^{t}))+\frac{Lm+1}{\rho}\sum_{i\in[m]}\left[\widehat{g}_{t,i}^{\ j\top}\widehat{q}_{t}^{\ j}-\alpha_{i}\right]^{+}\right)\leq\frac{(L\rho+L^{2}m^{2}+Lm)}{\rho}$ and $\sum_{t\in[T]}\sum_{j\in[M]}w_{t,j}^{2}\ell_{t,j}^{2}\leq\frac{TM(L\rho+L^{2}m^{2}+Lm)^{2}}{\rho^{2}}$. Thus, with probability at least $1-2\delta$, it holds:
@@
$\leq\mathcal{O}\left(\frac{M\ln T}{\eta}+\eta~{}{m^{4}L^{4}TM}+\gamma TLM+L\sqrt{T\ln\left(\frac{1}{\delta}\right)}+\frac{L}{\gamma}\ln\left(\frac{1}{\delta}\right)\right)+\frac{Lm+1}{\rho}\widehat{V}_{T,j^{*}},$ (20)
@@
$V_{T}-\widehat{V}_{T}=\mathcal{O}\left(mL|X||A|\sqrt{T\ln\left(\frac{mT|X||A|}{\delta}\right)}+m\ln(T)|X||A|C+\gamma TLM\right).$
@@
$\geq\sum_{t\in[T]}\sum_{j\in[M]}\frac{w_{t,j}\mathbb{I}(j_{t}=j)}{w_{t,j}+\gamma}\sum_{i\in[m]}\left(\left[g^{\circ}{}^{\top}q_{t}^{j}-\alpha_{i}\right]^{+}-\widehat{\xi}_{t,j}^{\ \top}q_{t}^{j}-\|q_{t}^{j}-\widehat{q}_{t}^{j}\|_{1}\right)$ (23b)
@@
$\mkern 80.0mu\cdot\sum_{i\in[m]}\left[(g^{\circ}-\mathbb{E}[g_{t,i}])^{\top}q_{t}^{j}\right]^{+}-\mathcal{O}\left(mL|X|\sqrt{|A|T\ln\left(\frac{T|X||A|}{\delta}\right)}\right),$ (23c)
@@
$=\mathcal{O}\left(m|X||A|\sqrt{T\ln\left(\frac{mMT|X||A|}{\delta}\right)}+m\ln T|X||A|C+L\sqrt{T\ln\frac{1}{\delta}}\right),$
@@
Upper-bound to $\sum_{t\in[T]}\sum_{j\in[M]}\frac{w_{t,j}\mathbb{I}(j_{t}=j)}{w_{t,j}+\gamma}\sum_{i\in[m]}\left[\left(g^{\circ}-\mathbb{E}[g_{t,i}]\right)^{\top}q_{t}^{j}\right]^{+}$.
@@
$\sum_{t\in[T]}\sum_{j\in[M]}\frac{w_{t,j}\mathbb{I}(j_{t}=j)}{w_{t,j}+\gamma}\sum_{i\in[m]}$ $\left[\left(g^{\circ}-\mathbb{E}[g_{t,i}]\right)^{\top}q_{t}^{j}\right]^{+}$
$\leq m\max_{i\in[m]}\sum_{t\in[T]}\sum_{j\in[M]}\mathbb{I}(j_{t}=j)\lVert g^{\circ}-\mathbb{E}[g_{t,i}]\rVert_{1}$
@@
$=\mathcal{O}\left(mL|X||A|\sqrt{T\ln\left(\frac{mMT|X||A|}{\delta}\right)}+m\ln(T)|X||A|C+\gamma TLM\right)$
@@
For any $\delta\in(0,1)$, with probability at least $1-14\delta$, it holds:
$\widehat{V}_{T,j^{*}}=\mathcal{O}\left(mL|X||A|\sqrt{T\ln\left(\frac{mMT|X||A|}{\delta}\right)}+m\ln(T)|X||A|C+L\frac{\ln\left(\frac{M}{\delta}\right)}{2\gamma}\right).$
@@
$\mkern 240.0mu+\left[g^{\circ}{}^{\top}q_{t}^{j^{*}}-\mathbb{E}[g_{t,i}]^{\top}q_{t}^{j^{*}}\right]^{+}+\lVert\widehat{q}_{t}^{j^{*}}-{q}_{t}^{j^{*}}\rVert_{1}\Bigg{)}$ (24b)
$\leq\sum_{t\in[T]}\frac{\mathbb{I}(j_{t}=j^{*})}{w_{t,j^{*}}+\gamma}\sum_{i\in[m]}\left(\left[\mathbb{E}[g_{t,i}]^{\top}q_{t}^{j^{*}}-\alpha_{i}\right]^{+}+\widehat{\xi}_{t,j^{*}}^{\ \top}q_{t}^{j^{*}}+\left[\left(g^{\circ}-\mathbb{E}[g_{t,i}]\right)^{\top}q_{t}^{j^{*}}\right]^{+}\right)$
@@
Upper-bound to $\sum_{t\in[T]}\frac{\mathbb{I}(j_{t}=j^{*})}{w_{t,j^{*}}+\gamma}\sum_{i\in[m]}\left[\left(g^{\circ}-\mathbb{E}[g_{t,i}]\right)^{\top}q_{t}^{j^{*}}\right]^{+}$.
@@
$\sum_{t\in[T]}\frac{\mathbb{I}(j_{t}=j^{*})}{w_{t,j^{*}}+\gamma}\sum_{i\in[m]}\left[\left(g^{\circ}-\mathbb{E}[g_{t,i}]\right)^{\top}q_{t}^{j^{*}}\right]^{+}\leq Lm\sqrt{2T\ln\left(\frac{1}{\delta}\right)}+mC.$
@@
We bound the quantity of interest as follows. With probability at least $1-10\delta$, it holds:
@@
$\leq\sum_{t\in[T]}\sum_{i\in[m]}\left[\mathbb{E}[g_{t,i}]^{\top}q_{t}^{j^{*}}-\alpha_{i}\right]^{+}+L\frac{\ln\left(\frac{M}{\delta}\right)}{2\gamma}$ (25a)
$\leq mV_{T,j^{*}}+L\frac{\ln(\frac{M}{\delta})}{2\gamma}$ (25b)
$=\mathcal{O}\left(m|X||A|L\sqrt{T\ln\left(\frac{mT|X||A|}{\delta}\right)}+m\ln(T)|X||A|C+L\frac{\ln(\frac{M}{\delta})}{2\gamma}\right),$ (25c)
where Inequality (25a) holds by Corollary 5 with probability at least $1-\delta$, and where $V_{T,j^{*}}$ in Inequality (25b) is the positive cumulative violation of costs constraints that Algorithm 2 would attain when instantiated with value of corruption $C=2^{j^{*}}$, if it were to run on its own, and thus Equality (25c) holds by Theorem 2 and Corollary 9 with probability at least $1-9\delta$ considering that, by definition of $j^{*}$ , $2^{j^{*}}$ is at most $2C$.
@@
$=\mathcal{O}\left(|X||A|\sqrt{T\ln\left(\frac{mMT|X||A|}{\delta}\right)}+\ln(T)|X||A|C+L\sqrt{T\ln\frac{1}{\delta}}\right),$

### v2

@@ F.1 Lagrangian formulation of the constrained optimization problem
$\min_{\lVert\lambda\rVert_{1}\in[0,L/\rho]}\max_{q\in\Delta(P)}\mathcal{Q}(\lambda,q)=\min_{\lambda\in\mathbb{R}^{m}_{\geq 0}}\max_{q\in\Delta(P)}\mathcal{Q}(\lambda,q),$
@@
$\min_{\lambda\in\mathbb{R}^{m}_{\geq 0}}\max_{q\in\Delta(P)}\mathcal{Q}(\lambda,q)$ $=\min\left\{\min_{\lVert\lambda\rVert_{1}\in[0,L/\rho]}\max_{q\in\Delta(P)}\mathcal{Q}(\lambda,q),\min_{\lVert\lambda\rVert_{1}\geq L/\rho}\max_{q\in\Delta(P)}\mathcal{Q}(\lambda,q)\right\}$
@@
$\text{OPT}_{r,G,\alpha}$ $=\max_{q\in\Delta(P)}\min_{\lambda\in\mathbb{R}^{m}_{\geq 0}}\mathcal{Q}(\lambda,q)$
@@ F.2 Preliminary results
F.2.1 Stability parameters
In the following sections, we will employ the stability parameters $\beta$ defined as follows:
  • $\beta_{1}=\mathcal{O}\left(L^{2}|X|^{2}|A|\ln\left(\frac{T|X||A|}{\delta}\right)\right)$
  • $\beta_{2}=\mathcal{O}\left(|X|^{2}|A|^{2}\log(T)\log\left(\nicefrac{{\log(T)}}{{\delta}}\right)\right)$
  • $\beta_{3}=\mathcal{O}\left(\ln(T)^{2}|X||A|\right)$
  • $\beta_{4}=\mathcal{O}\left(L^{2}|X|^{2}|A|\ln\left(\frac{mT|X||A|}{\delta}\right)\right)$
  • $\beta_{5}=\mathcal{O}\left(|X|^{2}|A|^{2}\log(T)\log\left(\nicefrac{{\log(T)}}{{\delta}}\right)\right)$
  • $\beta_{6}=\mathcal{O}\left(\ln(T)^{2}|X||A|\right)$
F.2.2 Omitted proofs and lemmas
@@
For any $\delta\in(0,1)$ and properly setting the learning rate $\eta$ such that $\eta\leq\frac{1}{2\Lambda m\left(\sqrt{\beta_{1}T}+\beta_{2}+\beta_{5}+\sqrt{\beta_{4}T}\right)}$, Algorithm 3 attains, with probability at least $1-2\delta$:
@@
$\hskip 28.45274pt+\left(\frac{m(mL+1)}{\rho}\beta_{5}+\beta_{2}\right)\nu_{T,j^{*}}+\left(\sqrt{\beta_{1}}+\left(\frac{m(Lm+1)}{\rho}\right)\sqrt{\beta_{4}}\right)\sqrt{T}\nu_{T,j*}$
$\leq\mathcal{O}\bigg{(}\frac{M\ln T}{\eta}+\eta~{}{m^{4}L^{4}TM}+\eta~{}M\ln(T)m^{4}L^{2}\beta_{5}^{2}+\eta~{}M\ln(T)\beta_{2}^{2}$
$\mkern 100.0mu+\eta T(\beta_{1}+L^{2}m^{4}\beta_{4})M\log(T)+\gamma TLM+L\sqrt{T\ln\left(\nicefrac{{1}}{{\delta}}\right)}+\frac{L}{\gamma}\ln\left(\nicefrac{{1}}{{\delta}}\right)\bigg{)}.$
@@
$\ell_{t,j}:=\frac{\mathbb{I}(j_{t}=j)}{w_{t,j}+\gamma}\left(\sum_{(x_{k}^{t},a_{k}^{t})}(1-r_{t}(x_{k}^{t},a_{k}^{t}))+\frac{Lm+1}{\rho}\sum_{i\in[m]}\left[\widehat{g}_{t,i}^{\ j\top}\widehat{q}_{t}^{\ j}-\alpha_{i}\right]^{+}\right),$
and $b_{t,j}$ for all $t\in[T]$, for all $j\in[M]$ as:
$b_{t,j}:=\left(\left(\frac{m(mL+1)}{\rho}\beta_{5}+\beta_{2}\right)+\left(\sqrt{\beta_{1}}+\frac{m(Lm+1)}{\rho}\sqrt{\beta_{4}}\right)\sqrt{T}\right)(\nu_{t,j}-\nu_{t-1,j}),$
with $\nu_{t,j}=\max_{\tau\in[t]}\frac{1}{w_{\tau,j}}$.
First we prove that $\eta w_{t,j}|\ell_{t,j}-b_{t,j}|\leq\nicefrac{{1}}{{2}}$ for all $t\in[T],j\in[M]$, to apply Lemma 17. It holds that $\eta w_{t,j}\lvert\ell_{t,j}\rvert\leq\frac{\eta(L\rho+L^{2}m^{2}+Lm)}{\rho}\leq\frac{1}{2}$ for all $j\in[M]$, for all $t\in[T]$ as long as $\eta\leq\frac{\rho}{2(L\rho+L^{2}m^{2}+Lm)}\leq\frac{\rho}{2(L^{2}m^{2}+Lm)}$, which is true if $\eta\leq\frac{\rho}{2Lm(Lm+1)}$. It also holds that
$\eta w_{t,j}|b_{t,j}|$ $=\eta w_{t,j}\left(\left(\frac{m(Lm+1)}{\rho}\beta_{5}+\beta_{2}\right)+\left(\frac{m(Lm+1)}{\rho}\sqrt{\beta_{4}}+\sqrt{\beta_{1}}\right)\sqrt{T}\right)(\nu_{t,j}-\nu_{t-1,j})$
$\leq\eta\left(\left(\frac{m(Lm+1)}{\rho}\beta_{5}+\beta_{2}\right)+\left(\frac{m(Lm+1)}{\rho}\sqrt{\beta_{4}}+\sqrt{\beta_{1}}\right)\sqrt{T}\right)\left(1-\frac{\nu_{t-1,j}}{\nu_{t,j}}\right)$
$\leq\eta\left(\left(\frac{m(Lm+1)}{\rho}\beta_{5}+\beta_{2}\right)+\left(\frac{m(Lm+1)}{\rho}\sqrt{\beta_{4}}+\sqrt{\beta_{1}}\right)\sqrt{T}\right)$
$\leq\frac{1}{2},$
if $\eta\leq\frac{1}{2\Lambda m\left(\sqrt{\beta_{1}T}+\beta_{2}+\beta_{5}+\sqrt{\beta_{4}T}\right)}$, where we used the fact that $\nu_{t,j}\neq\nu_{t-1,j}\iff\nicefrac{{1}}{{w_{t,j}}}=\nu_{t,j}$. Thus, if the previous conditions on $\eta$ hold, and notice that the second condition implies the first, Algorithm 3 attains, by Lemma 17 :
@@
$\leq\frac{M\ln T}{\eta}+2\eta\frac{TM(L\rho+L^{2}m^{2}+Lm)^{2}}{\rho^{2}}$
$\ \ \ \ +2\eta\left(2\left(\frac{m(mL+1)}{\rho}\beta_{5}+\beta_{2}\right)^{2}M\ln\left(T\right)+2T\left(\sqrt{\beta_{1}}+\left(\frac{m(Lm+1)}{\rho}\right)\sqrt{\beta_{4}}\right)^{2}M\ln(T)\right)$
$\mkern 250.0mu+\frac{Lm+1}{\rho}\widehat{V}_{T,j^{*}}+\sum_{t\in[T]}\sum_{j\in[M]}w_{t,j}b_{t,j}-\sum_{t\in[T]}b_{t,j^{*}},$ (19)
where we used the following inequalities:
  • First inequality:
    $\sum_{t\in[T]}\sum_{j\in[M]}w_{t,j}^{2}(\ell_{t,j}-b_{t,j})^{2}\leq 2\sum_{t\in[T]}\sum_{j\in[M]}w_{t,j}^{2}\ell_{t,j}^{2}+2\sum_{t\in[T]}\sum_{j\in[M]}w_{t,j}^{2}b_{t,j}^{2},$
  • Second inequality:
    $\left(\sum_{(x_{k}^{t},a_{k}^{t})}(1-r_{t}(x_{k}^{t},a_{k}^{t}))+\frac{Lm+1}{\rho}\sum_{i\in[m]}\left[\widehat{g}_{t,i}^{\ j\top}\widehat{q}_{t}^{\ j}-\alpha_{i}\right]^{+}\right)\leq\frac{(L\rho+L^{2}m^{2}+Lm)}{\rho},$
  • Third inequality:
    $\sum_{t\in[T]}\sum_{j\in[M]}w_{t,j}^{2}\ell_{t,j}^{2}\leq\frac{TM(L\rho+L^{2}m^{2}+Lm)^{2}}{\rho^{2}},$
and that, it holds:
$\sum_{t\in[T]}\sum_{j\in[M]}w_{t,j}^{2}b_{t,j}^{2}$
$=\sum_{t\in[T]}\sum_{j\in[M]}(w_{t,j}b_{t,j})^{2}$
$\leq\left(\left(\frac{m(Lm+1)}{\rho}\beta_{5}+\beta_{2}\right)+\left(\frac{m(Lm+1)}{\rho}\sqrt{\beta_{4}}+\sqrt{\beta_{1}}\right)\sqrt{T}\right)^{2}\sum_{j\in[M]}\sum_{t\in[T]}\left(\frac{1}{\nu_{t,j}}(\nu_{t,j}-\nu_{t-1,j})\right)^{2}$ (20a)
$\leq\left(2\left(\frac{m(mL+1)}{\rho}\beta_{5}+\beta_{2}\right)^{2}+2T\left(\frac{m(Lm+1)}{\rho}\sqrt{\beta_{4}}+\sqrt{\beta_{1}}\right)^{2}\right)\sum_{j\in[M]}\sum_{t\in[T]}\left(1-\frac{\nu_{t-1,j}}{\nu_{t,j}}\right)^{2}$
$\leq\left(2\left(\frac{m(mL+1)}{\rho}\beta_{5}+\beta_{2}\right)^{2}+2T\left(\frac{m(Lm+1)}{\rho}\sqrt{\beta_{4}}+\sqrt{\beta_{1}}\right)^{2}\right)\sum_{j\in[M]}\sum_{t\in[T]}\left(1-\frac{\nu_{t-1,j}}{\nu_{t,j}}\right)$
$\leq\left(2\left(\frac{m(mL+1)}{\rho}\beta_{5}+\beta_{2}\right)^{2}+2T\left(\frac{m(Lm+1)}{\rho}\sqrt{\beta_{4}}+\sqrt{\beta_{1}}\right)^{2}\right)\sum_{j\in[M]}\sum_{t\in[T]}\ln\left(\frac{\nu_{t,j}}{\nu_{t-1,j}}\right)$ (20b)
$\leq\left(2\left(\frac{m(mL+1)}{\rho}\beta_{5}+\beta_{2}\right)^{2}+2T\left(\frac{m(Lm+1)}{\rho}\sqrt{\beta_{4}}+\sqrt{\beta_{1}}\right)^{2}\right)\sum_{j\in[M]}\ln\left(\prod_{t\in[T]}\frac{\nu_{t,j}}{\nu_{t-1,j}}\right)$
$\leq\left(2\left(\frac{m(mL+1)}{\rho}\beta_{5}+\beta_{2}\right)^{2}+2T\left(\frac{m(Lm+1)}{\rho}\sqrt{\beta_{4}}+\sqrt{\beta_{1}}\right)^{2}\right)\sum_{j\in[M]}\ln\left(\frac{\nu_{T,j}}{\nu_{0,j}}\right)$
$\leq\left(2\left(\frac{m(mL+1)}{\rho}\beta_{5}+\beta_{2}\right)^{2}+2T\left(\frac{m(Lm+1)}{\rho}\sqrt{\beta_{4}}+\sqrt{\beta_{1}}\right)^{2}\right)M\ln\left(T\right),$ (20c)
where Inequality (20a) is true since $\nu_{t,j}-\nu_{t-1,j}\neq 0$ only when $w_{t,j}=\nicefrac{{1}}{{\nu_{t,j}}}$ by definition, Inequality (20b) holds since $1-a\leq-\ln a$, and Inequality (20c) holds since by definition $\nu_{T,j}\leq T$ and $\nu_{0,j}=M$. Notice also that, following a similar reasoning, it holds:
$\sum_{t\in[T]}w_{t,j}b_{t,j}-\sum_{t\in[T]}b_{t,j^{*}}$
$=\left(\left(\frac{m(Lm+1)}{\rho}\beta_{5}+\beta_{2}\right)+\left(\frac{m(Lm+1)}{\rho}\sqrt{\beta_{4}}+\sqrt{\beta_{1}}\right)\sqrt{T}\right)\sum_{t\in[T]}\sum_{j\in[M]}\left(1-\frac{\nu_{t-1,i}}{\nu_{t,i}}\right)$
$\quad-\left(\left(\frac{m(Lm+1)}{\rho}\beta_{5}+\beta_{2}\right)+\left(\frac{m(Lm+1)}{\rho}\sqrt{\beta_{4}}+\sqrt{\beta_{1}}\right)\sqrt{T}\right)\sum_{t\in[T]}(\nu_{t,j^{*}}-\nu_{t-1,j^{*}})$
$\leq\mathcal{O}\left(m^{2}L\beta_{5}M\log(T)+\beta_{2}M\log(T)+(\sqrt{\beta_{1}}+Lm^{2}\sqrt{\beta_{4}})\sqrt{T}M\log(T)\right)$
$\quad-\left(\left(\frac{m(Lm+1)}{\rho}\beta_{5}+\beta_{2}\right)+\left(\frac{m(Lm+1)}{\rho}\sqrt{\beta_{4}}+\sqrt{\beta_{1}}\right)\sqrt{T}\right)\nu_{T,j^{*}}$
Thus, with probability at least $1-2\delta$, it holds:
@@
$\leq\mathcal{O}\bigg{(}\frac{M\ln T}{\eta}+\eta~{}{m^{4}L^{4}TM}+\eta~{}M\ln(T)m^{4}L^{2}\beta_{5}^{2}+\eta~{}M\ln(T)\beta_{2}^{2}$
$\mkern 70.0mu+\eta T(\beta_{1}+L^{2}m^{4}\beta_{4})M\log(T)+\gamma TLM+L\sqrt{T\ln\left(\nicefrac{{1}}{{\delta}}\right)}+\frac{L}{\gamma}\ln\left(\nicefrac{{1}}{{\delta}}\right)\bigg{)}+\frac{Lm+1}{\rho}\widehat{V}_{T,j^{*}}$
$\mkern 70.0mu-\left(\frac{m(mL+1)}{\rho}\beta_{5}+\beta_{2}\right)\nu_{T,j^{*}}-\left(\sqrt{\beta_{1}}+\left(\frac{m(Lm+1)}{\rho}\right)\sqrt{\beta_{4}}\right)\sqrt{T}\nu_{T,j*},$ (22)
@@
$V_{T}-\widehat{V}_{T}=\mathcal{O}\left(mL|X|\sqrt{|A|T\ln\left(\frac{mT|X||A|}{\delta}\right)}+m\ln(T)|X||A|C+\gamma TLM\right).$
@@
$\geq\sum_{t\in[T]}\sum_{j\in[M]}\frac{w_{t,j}\mathbb{I}(j_{t}=j)}{w_{t,j}+\gamma}\sum_{i\in[m]}\left(\left[g_{i}^{\circ\top}q_{t}^{j}-\alpha_{i}\right]^{+}-\widehat{\xi}_{t,j}^{\ \top}q_{t}^{j}-\|q_{t}^{j}-\widehat{q}_{t}^{j}\|_{1}\right)$ (25b)
@@
$\mkern 80.0mu\cdot\sum_{i\in[m]}\left[(g^{\circ}_{i}-\mathbb{E}[g_{t,i}])^{\top}q_{t}^{j}\right]^{+}-\mathcal{O}\left(mL|X|\sqrt{|A|T\ln\left(\frac{T|X||A|}{\delta}\right)}\right),$ (25c)
@@
$=\mathcal{O}\left(m\sqrt{|X||A|LT\ln\left(\frac{mMT|X||A|}{\delta}\right)}+m\ln T|X||A|C+L\sqrt{T\ln\frac{1}{\delta}}\right),$
@@
Upper-bound to $\sum_{t\in[T]}\sum_{j\in[M]}\frac{w_{t,j}\mathbb{I}(j_{t}=j)}{w_{t,j}+\gamma}\sum_{i\in[m]}\left[\left(g^{\circ}_{i}-\mathbb{E}[g_{t,i}]\right)^{\top}q_{t}^{j}\right]^{+}$.
@@
$\sum_{t\in[T]}\sum_{j\in[M]}\frac{w_{t,j}\mathbb{I}(j_{t}=j)}{w_{t,j}+\gamma}\sum_{i\in[m]}$ $\left[\left(g_{i}^{\circ}-\mathbb{E}[g_{t,i}]\right)^{\top}q_{t}^{j}\right]^{+}$
$\leq m\max_{i\in[m]}\sum_{t\in[T]}\sum_{j\in[M]}\mathbb{I}(j_{t}=j)\lVert g_{i}^{\circ}-\mathbb{E}[g_{t,i}]\rVert_{1}$
@@
$=\mathcal{O}\left(mL|X|\sqrt{|A|T\ln\left(\frac{mMT|X||A|}{\delta}\right)}+m\ln(T)|X||A|C+\gamma TLM\right)$
@@
For any $\delta\in(0,1)$, with probability at least $1-16\delta$, it holds:
$\widehat{V}_{T,j^{*}}$ $\leq\mathcal{O}\left(mL|X|\sqrt{|A|T\ln\left(\frac{mMT|X||A|}{\delta}\right)}+m\beta_{6}C+m\ln(T)|X||A|C+Lm\frac{\ln\left(\frac{M}{\delta}\right)}{2\gamma}\right)$
$\mkern 350.0mu+m\sqrt{\beta_{4}T}\nu_{T,j^{*}}+m\beta_{5}\nu_{T,j^{*}}.$
@@
$\mkern 240.0mu+\left[g^{\circ}_{i}{}^{\top}q_{t}^{j^{*}}-\mathbb{E}[g_{t,i}]^{\top}q_{t}^{j^{*}}\right]^{+}+\lVert\widehat{q}_{t}^{j^{*}}-{q}_{t}^{j^{*}}\rVert_{1}\Bigg{)}$ (26b)
$\leq\sum_{t\in[T]}\frac{\mathbb{I}(j_{t}=j^{*})}{w_{t,j^{*}}+\gamma}\sum_{i\in[m]}\left(\left[\mathbb{E}[g_{t,i}]^{\top}q_{t}^{j^{*}}-\alpha_{i}\right]^{+}+\widehat{\xi}_{t,j^{*}}^{\ \top}q_{t}^{j^{*}}+\left[\left(g^{\circ}_{i}-\mathbb{E}[g_{t,i}]\right)^{\top}q_{t}^{j^{*}}\right]^{+}\right)$
@@
Upper-bound to $\sum_{t\in[T]}\frac{\mathbb{I}(j_{t}=j^{*})}{w_{t,j^{*}}+\gamma}\sum_{i\in[m]}\left[\left(g^{\circ}_{i}-\mathbb{E}[g_{t,i}]\right)^{\top}q_{t}^{j^{*}}\right]^{+}$.
@@
$\sum_{t\in[T]}\frac{\mathbb{I}(j_{t}=j^{*})}{w_{t,j^{*}}+\gamma}\sum_{i\in[m]}\left[\left(g^{\circ}_{i}-\mathbb{E}[g_{t,i}]\right)^{\top}q_{t}^{j^{*}}\right]^{+}\leq Lm\sqrt{2T\ln\left(\frac{1}{\delta}\right)}+mC.$
@@
We bound the quantity of interest as follows. With probability at least $1-11\delta$, it holds:
@@
$\leq m\sqrt{\beta_{4}T}\nu_{T,j^{*}}+m\beta_{5}\nu_{T,j^{*}}+2m\beta_{6}C+Lm\frac{\ln\left(\frac{M}{\delta}\right)}{2\gamma},$ (27a)
thank to Corollary 5 and Corollary 6 .
@@
$=\mathcal{O}\left(\sqrt{|X||A|LT\ln\left(\frac{mMT|X||A|}{\delta}\right)}+\ln(T)|X||A|C+L\sqrt{T\ln\frac{1}{\delta}}\right),$